OpenStudy (anonymous):

Easton simplified the following expression: (x2y6z5)(x4y5z3) If he writes his answer in the form of xaybzc, what is the value of b, the exponent on y? Numerical Answers Expected!

OpenStudy (anonymous):

just add the exponents of y you already have remember this a^n*a^m=a^n+m (x^2y^6z^5)(x^4y^5z^3)=(x^2x^4)(y^6y^5)(z^5z^3) Example: T^4*T^7=T^4+7=T^11 You add the EXPONENTS, not the variables.